3) Evaluate. 81 ÷ 9 – 16 ÷ 8
i thought was 8 but i was wrong.

Answer by fractalier(2101)

Order of operations tells us that division is done before subtraction no matter how they are listed in the problem, so we have
81 ÷ 9 – 16 ÷ 8 =
9 - 2 =
7